Find a polynomial function of degree 3 with real coefficients that has the given zeros.
-1, 2, -4

Answer by ikleyn(52788)

.
The simplest polynomial of degree 3 with these zeroes is

    p(x) = (x-(-1))*(x-2)*(x-(-4)) = (x+1)*(x-2)*(x+4).




A GENERAL polynomial of degree 3 with these zeroes is

    P(x) = a*(x+1)*(x-2)*(x+4),

where "a" is any real coefficient (different from 0).